# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v1

@@ 2.4. Emerging Research Tasks
In Section 2.1, we mentioned one of the inherent drawbacks of WsVAD, which violates the open-set problem property of the AD task by only detecting specific anomalies predefined in the training set and not being able to cope with diverse and arbitrary anomalous events in the open world. Action Recognition (AR) (Kong and Fu, 2022; Sun et al., 2022) and VAD aim to understand specific behaviors. They can learn from each other’s research ideas in data modality and feature learning, and the pre-trained AR models on large-scale video datasets are expected to be directly used for feature extraction and anomaly semantic cue mining in VAD. However, AR follows the closed-set task setting, which can only model the spatial-temporal patterns of known categories of behaviors but cannot empower the model to recognize out-of-distribution samples. In response, researchers have proposed various open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022) schemes to break through the above barriers, as shown in Fig. 7(b).
@@
For example, OpenVAD in (Zhu et al., 2022) aims to integrate the advantages of UVAD, which can handle arbitrary anomalous events, and WAED, which has a low false alarm rate in detecting specific anomalies. The proposed method integrates deep evidence learning and normalized flow into the MIL to equip WsVAD with the ability to identify unknown anomalies by quantifying uncertainty. Acsintoae et al. (Acsintoae et al., 2022) propose a dataset for supervised OSVAD, named UBnormal, that maintains the task’s open-set properties. Since the anomalous behavior in this dataset is generated through the VAD engine, it comes with fine pixel-level labels, making it possible to train VAD models in supervised learning. In short, this dataset attempts to bridge closed supervised learning and open anomaly detection, and experiments show that it can improve performance without compromising the open-set properties of existing VAD models. OVVAD is closest in setting to the OSR task and aims to learn a multi-classifier capable of detecting and classifying all known and unknown anomalous events. Compared to OSVAD, OVVAD is more in line with the display requirements of scenarios such as autonomous driving. The first OVVAD model proposed by Wu et al. (Wu et al., 2024b) splits the task into two complementary tasks, i.e., AD and anomaly classification, and jointly optimizes them using the knowledge from the large models.
2.5. Applications
IoVT (Chen, 2020) has been widely used in daily life so that VAD models can be directly deployed in off-the-shelf cameras without additional hardware costs to realize contactless real-time environmental monitoring of factories, roads, households, and public places. In addition, 5G communication technology has driven the rapid development of mobile internet, and the number of short videos uploaded to online video-sharing platforms from smartphone terminals alone exceeds 400 million per day, making automation of content regulation imminent. We expect that this article will not only inspire VAD researchers to conduct academic studies but also attract the attention of practitioners in the transportation management, modern industrial, retail commerce, and public service industries.
The application scenarios of NSVAD include: a) Traffic accident detection and road security monitoring in smart transportation. Such problems have been the focus of WsVAD. The first weakly supervised dataset, UCF-Crime (Sultani et al., 2018), contains such traffic anomalies. The subsequently proposed ROADMAP dataset (Wang et al., 2021) is all captured from real traffic scenarios, which creates a benchmark for video-driven spatial-temporal anomaly detection research in open space. b) Accident discovery and production line management in model industry, etc. Video can collect large-scale spatial-temporal information without contact, and has the advantages of high information density, wide perception field of view, stable data quality, and adaptability to extreme environments compared to sensors such as temperature and humidity. Therefore, the use of video cameras to monitor the state of industrial production has gained the attention of researchers. As an example, we propose a multi-branch normality learning framework in (Liu et al., 2024b) to cope with high-motion dynamics and complex appearance patterns in industrial scenes. Experiments conducted on realistic datasets validate the effectiveness of the proposed AMP-Net, which reflects the potential of VAD in industrial scenarios. c) Commercial entities directly facing people, such as retail, often face the risk of economic losses due to theft, robbery, and fraudulent behaviors. Both the UCF-crime (Sultani et al., 2018) and XD-violence (Wu et al., 2020) datasets capture such scenarios of violations, providing benchmarks for building crime detection systems for public commercial areas. In addition, VAD can provide new ideas for customer behavior analysis, i.e., detecting abnormal customer flows to guide the adjustment of store layout and business strategies, etc. d) The public service usually needs to bear data privacy concerns, especially for surveillance videos containing biological information (e.g., faces, body behaviors, and sounds). Human inspection is not only time-consuming and laborious but also may cause sensitive information leakage. In contrast, VAD can analyze video on the user side and transfer only the segments of concern to the third parties.
@@ 3. Unsupervised Video Anomaly Detection
UVAD follows the general setup of semantic anomaly detection tasks, where only easily collectible regular instances are used to train models describing the normality of videos, detecting anomalies by measuring deviations between test samples and the learned model. From traditional machine learning to deep representation learning, UVAD has undergone multiple revolutions in feature extraction and normality learning, leading VAD to become a key issue in the AD and CV communities. With its close ties to the AD community and its long-standing development history, UVAD has long been regarded by researchers as the mainstream research route of NSVAD algorithms. As a result, existing surveys (Nayak et al., 2021; Ramachandra et al., 2020) typically focus on reviewing UVAD literature, lacking in-depth discussion of emerging weakly supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Zaheer et al., 2022; Pang et al., 2020) routes, and overlooking novel tasks such as open-set and open-word detection. This article not only systematically reviews the latest developments in these new routes and tasks but also provides foundational knowledge and timely and classic methods of UVAD in this section.
@@
Based on the means of normality learning and deviation calculation principles, existing deep learning methods are generally divided into three categories: distance-based (Saligrama and Chen, 2012; Dalal et al., 2006; Dalal and Triggs, 2005), probability-based (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), and reconstruction-based (Liu et al., 2024b, 2018b; Bao et al., 2022; Luo et al., 2017). From our perspective, distance-based UVAD methods are a more general form of probability-based and reconstruction-based methods because probability deviation and reconstruction error calculation are fundamentally just different distance measurement methods. Specifically, distance-based methods include using single-classifier learning of video sequence spatial-temporal representations in the deep feature space, such as OC-SVM and OC-NN (Ionescu et al., 2019; Sun et al., 2019). The drawback of such methods is that the trained models cannot be incrementally compatible with new data, leading to the need to retrain classifiers from scratch when new data is generated, such as in scenarios involving scene transitions. Another type of distance-based method is to use Gaussian mixture models (Wu et al., 2010) to model video feature normal vectors and measure deviations using Mahalanobis distance. In contrast, probability-based methods attempt to map the spatial-temporal representations of regular events into a probability framework and discriminate anomaly instances by measuring differences in probability distributions. Such methods tend to use traditional models such as Markov random fields (Benezeth et al., 2009) to build probability space. Deep learning-based attempts have encountered significant increases in computational costs and are noticeably slower in inference speeds.
@@
In fact, the most prevailing deep UVAD approach is reconstruction-based (Ramachandra et al., 2020). On the one hand, reconstruction-based methods (Liu et al., 2018b; Park et al., 2020; Liu et al., 2024b) aim to train models to represent the general spatial-temporal patterns of regular events through self-supervised proxy tasks, benefiting from advances in video self-supervised learning (Jing and Tian, 2020) and deep neural networks. On the other hand, such methods avoid complex mathematical computations, which are easy to implement and exhibit excellent performance, thus being widely praised by existing researchers. The premise of reconstruction-based methods is that generative models trained on massive normal samples can effectively infer the spatial-temporal patterns of regular events. For anomalous events, the performance of proxy tasks will significantly decrease, and the resulting error can be used as a quantitative basis for measuring deviations to calculate anomaly scores. Common generative models include deep autoencoders (Hasan et al., 2016; Chong and Tay, 2017; Liu et al., 2022a, c; Park et al., 2020), variational autoencoders (Fan et al., 2020; Yan et al., 2018), and generative adversarial networks (Chen et al., 2021; Huang et al., 2021; Nguyen et al., 2020; Li et al., 2022; Cai et al., 2021a; Zhang et al., 2022). Proxy tasks include reconstructing input sequences and predicting future frames, which belong to pixel-level image generation. However, most reconstruction-based methods only calculate frame-level errors as anomaly scores without performing spatial localization, as the spatial contribution of anomalies is typically not significant. Spatial localization can only serve as a quantitative visualization results.
@@ 3.1. Taxonomy and Advances
This article also systematically reviews deep learning-driven UVAD methods, providing a taxonomy that can aid in understanding the current state of research and inspire further exploration. However, due to space constraints and the emphasis of the tutorial paper on guiding beginners, we only present the underlying logic of the proposed taxonomy and summarize the research trends of the latest advances in the main text message. For more detailed description of the existing UVAD methods, please refer to Section 1.1 of Supplement.
@@
The methods based on distance and probability described earlier focus more on handcrafted features (Cheng et al., 2015; Ravanbakhsh et al., 2018; Shi et al., 2021; Ravanbakhsh et al., 2017) and traditional classification (Tran and Hogg, 2017; Turchini et al., 2017; Vu et al., 2017) models, which have been primarily surpassed by distance-based methods in the era of deep learning. Therefore, existing classification systems appear outdated in delineating recent advancements and reflecting research trends, failing to highlight the latest challenges and directions. To address this, inspired by data preprocessing techniques and forms of deep learning modeling, we categorize UVAD into two main classes: Global Normality Learning (GNL) and Local Prototype Modeling (LPM), as presented in Table. 2.
@@
GNL utilizes the entire video sequence as input, often requiring no additional preprocessing such as spatial-temporal cube partitioning or foreground object extraction, and employs end-to-end deep neural networks to directly learn video ecologies (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Chang et al., 2020; Liu et al., 2024b, 2023c). Over the years, researchers have believed that videos possess two informational dimensions, namely spatial and temporal, corresponding to appearance and motion, requiring different approaches to capture their normality. Therefore, within our UVAD taxonomy, GNL is further subdivided into Single-Proxy Task (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020) and Multi-Proxy Task methods (Chang et al., 2020; Liu et al., 2022c, 2024b). In contrast, LPM methods argue that video data contain a plethora of redundant information tantamount to clues related to normality. Thus, they opt to use spatial-temporal cubes (Mahadevan et al., 2010; Roshtkhari and Levine, 2013; Sabokrou et al., 2017; Liu et al., 2021a) or foreground objects (Bao et al., 2022; Liu et al., 2023a) containing dense and effective information as network inputs instead of the entire video sequence, focusing on prototype feature learning of local patch. We classify such methods into spatial-temporal Patch-based and Object-driven methods.
@@ 3.1.1. Global Normality Learning
Convolutional neural network-driven deep representation models can directly learn task-relevant spatial-temporal representations from raw video sequences and can adapt to different scale inputs and feature dimension requirements through simple structural adjustments. Global normality learning aims to learn video normality directly from complete RGB videos or optical flow sequences. Compared to local prototype modeling, GNL requires no additional data preparation and is easy to optimize.
The earliest methods did not distinguish between spatial and temporal information, typically using only RGB videos as input and employing a single self-supervised proxy task (e.g., sequence reconstruction or future frame prediction) to broadly learn spatial-temporal normality. Generally, methods based on a single proxy task focus on designing more efficient single-stream end-to-end deep structures. Recent efforts include the introduction of 3D convolutional networks (Gong et al., 2019) and convolutional long short-term memory networks (Luo et al., 2017) to enhance the representation capability of spatial-temporal features. Subsequent researchers discovered that spatial and temporal information have different characteristics, with the former focusing on local pixel inference and the latter on modeling global dynamics. Moreover, the addition of proxy tasks such as reconstruction and prediction losses often brings additional performance gains without significantly increasing training costs. Thus, they proposed introducing additional proxy tasks within the GNL framework. In addition to the separation of spatial-temporal normality learning (Liu et al., 2022c; Chang et al., 2021) and simple proxy task stacking (Zhao et al., 2017; Chen et al., 2020), recent work has also explored novel tasks such as appearance-motion consistency (Cai et al., 2021b), spatial-temporal coherence (Cheng et al., 2023; Ning et al., 2024), and correlation (Zhao et al., 2022).

### v2

@@ 2.4. Emerging Research Tasks
In Section 2.1, we mentioned one of the inherent drawbacks of WsVAD, which violates the open-set problem property of the AD task by only detecting specific anomalies predefined in the training set and not being able to cope with diverse and arbitrary anomalous events in the open world. Action Recognition (AR) (Kong and Fu, 2022; Sun et al., 2022) and VAD aim to understand specific behaviors. They can learn from each other’s research ideas in data modality and feature learning, and the pre-trained AR models on large-scale video datasets are expected to be directly used for feature extraction and anomaly semantic cue mining in VAD. However, AR follows the closed-set task setting, which can only model the spatial-temporal patterns of known categories of behaviors but cannot empower the model to recognize out-of-distribution samples. In response, researchers have proposed various open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022) schemes to break through the above barriers, as shown in Fig. 6(b). Zhang et al. (Zhang et al., 2024b) introduced the concept of glance annotation, where a single frame from an anomalous event is randomly labeled and used as an enhanced supervision signal for training weakly supervised VAD models. They provided glance annotations for UCF-Crime and XD-Violence datasets, achieving a 5% improvement in frame-level AUC compared to the state-of-the-art, demonstrating this setting’s outstanding potential for balancing annotation costs and model performance.
@@
OpenVAD in (Zhu et al., 2022) aims to integrate the advantages of UVAD, which can handle arbitrary anomalous events, and WAED, which has a low false alarm rate in detecting specific anomalies. The proposed method integrates evidential deep learning and normalized flow into the MIL to equip WsVAD with the ability to identify unknown anomalies by quantifying uncertainty. Acsintoae et al. (Acsintoae et al., 2022) propose a dataset for supervised OSVAD, named UBnormal, that maintains the task’s open-set properties. Since the anomalous behavior in this dataset is generated through the VAD engine, it comes with fine pixel-level labels, making it possible to train VAD models in supervised learning. In short, this dataset attempts to bridge closed supervised learning and open anomaly detection, and experiments show that it can improve performance without compromising the open-set properties of existing VAD models. OVVAD is closest in setting to the OSR task and aims to learn a multi-classifier capable of detecting and classifying all known and unknown anomalous events. Compared to OSVAD, OVVAD is more in line with the display requirements of scenarios such as autonomous driving. The first OVVAD model proposed by Wu et al. (Wu et al., 2024c) splits the task into two complementary tasks, i.e., AD and anomaly classification, and jointly optimizes them using the knowledge from the large models.
2.5. Language model-based VAD
Language models like Generative Pre-Training (GPT) (Radford, 2018) and Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021b) exhibit outstanding zero-shot learning and multimodal information processing abilities, showing great potential in VAD. Language models (Yang et al., 2024c; Teng et al., 2024) can provide accurate textual descriptions of video frames, enhancing the semantic mining capability of existing vision-based VAD models. For instance, Chen et al. (Chen et al., 2023) used a language model-based captioning network to obtain textual descriptions of video sequences, which, after being embedded in a text embedding network, were fused with visual features as inputs to the anomaly detector. Their proposed Text Empowered Video Anomaly Detection (TEVAD) efficiently captures abstract semantics of anomaly events and enhances the interpretability of VAD models. Pu et al. (Pu et al., 2024) introduced a Prompt-Enhanced Learning (PEL) module, using knowledge-based prompts to incorporate semantic priors, improving the discriminative power of visual features in weakly supervised VAD, while ensuring separability between anomalous subcategories. Wu et al. (Wu et al., 2024b) proposed VAR, which efficiently detects specific anomalies based on cross-modal learning (e.g., language descriptions and synchronized audio). They designed an Anomaly-Led Alignment Network (ALAN) using BERT (Kenton and Toutanova, 2019) to process text information and incorporate a pretext task to enhance semantic alignment between video-text fine-grained representations.
In addition, visual-language associations can serve as effective cues for detecting video anomalies, and the pre-learned visual-text consistency in large Visual-Language Models (VLMs) can be efficiently transferred to VAD. Kim et al. (Kim et al., 2023) used large language models to generate textual descriptions of video frames and detected anomalous frames by calculating the cosine similarity between input frames and their textual descriptions using CLIP. Text Prompt with Normality Guidance (Yang et al., 2024b) leverages the language-visual knowledge of the CLIP model to align video frames with textual descriptions of events, generating more accurate pseudo-labels for WsVAD, thus improving model performance. Inspired by CLIP, Wu et al. (Wu et al., 2024e) proposed a dual-branch structure called VadCLIP that utilizes the connection between vision and language for weakly supervised VAD. One branch uses only visual features for coarse-grained binary classification, while the other fully exploits fine-grained visual-language alignment to transfer CLIP’s knowledge to VAD. They also used a pre-trained VLM to learn spatio-temporal prompt embeddings for weakly supervised VAD and localization. The proposed STPrompt (Wu et al., 2024d) leverages the knowledge learned from VLMs and integrates natural motion priors from the original video to learn prompt embeddings consistent with spatial-temporal regions of the video (e.g., patches of individual frames) to identify specific local regions of anomalies.
Research has shown that multimodal LLMs can learn prototype patterns of normal events in a zero-shot manner and describe any anomalies in open-set settings, significantly improving the generality and adaptability of VAD models. Zanella et al. (Zanella et al., 2024) proposed Language-based VAD (LAVAD), an unsupervised learning paradigm that leverages pre-trained LLMs and existing VLMs to train video anomaly detectors. They used VLMs to generate textual descriptions of video frames and designed a prompt mechanism to unlock LLMs’ potential in temporal aggregation and anomaly score estimation, enabling direct VAD execution. Lv et al. (Lv and Sun, 2024) introduced Video-LLaMA into VAD, aiming to break threshold limitations and improve model interpretability. The authors proposed a three-stage training method to improve VLLM’s training efficiency. In the AnomalyRuler, during the inductive stage, a small set of normal reference videos was provided to the LLM, enabling it to summarize normal patterns to induce rules for anomaly detection. In the Deduction stage, these induced rules were applied to detect anomalous frames in test videos. Hawk in (Tang et al., 2024) uses an interactive VLM to accurately interpret video anomalies, answering VAD-related questions. It constructs an auxiliary consistency loss within the motion and video space, guiding the video branch to focus on motion modalities and establishing explicit supervision between actions and language to improve the accuracy of interpretation. Zhang et al. (Zhang et al., 2024c) developed a large-scale multimodal VAD instruction tuning benchmark called VAD-Instruct50k, used to build unbiased and interpretable VAD systems, as well as Holmes-VAD for anomaly event localization and interpretation.
@@ 3. Unsupervised Video Anomaly Detection
UVAD follows the general setup of semantic anomaly detection tasks, where only easily collectible regular instances are used to train models describing the normality of videos, detecting anomalies by measuring deviations between test samples and the learned model (Cheng et al., 2024a, b; Wang et al., 2021b). From traditional machine learning to deep representation learning (Yang et al., 2024a; Wang et al., 2024b; Yang et al., 2023b), UVAD has undergone multiple advancements in feature extraction and normality learning, leading VAD to become a key issue in the AD and CV communities. With its close ties to the AD community and its long-standing development history, UVAD has long been regarded by researchers as the mainstream research route of NSVAD algorithms. As a result, existing surveys (Nayak et al., 2021; Ramachandra et al., 2020) typically focus on reviewing UVAD literature, lacking in-depth discussion of emerging weakly supervised (Feng et al., 2021a; Tian et al., 2021) and fully unsupervised (Zaheer et al., 2022; Pang et al., 2020) routes, and overlooking novel tasks such as open-set and open-word detection. This article not only systematically reviews the latest developments in these new routes and tasks but also provides foundational knowledge and timely and classic methods of UVAD in this section.
@@
Based on the means of normality learning and deviation calculation principles, existing deep learning methods are generally divided into three categories: distance-based (Saligrama and Chen, 2012; Dalal et al., 2006; Dalal and Triggs, 2005), probability-based (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), and reconstruction-based (Liu et al., 2024b, 2023f; Bao et al., 2022; Luo et al., 2017). From our perspective, distance-based UVAD methods are a more general form of probability-based and reconstruction-based methods because probability deviation and reconstruction error calculation are fundamentally just different distance measurement methods. Specifically, distance-based methods include using single-classifier learning of video sequence spatial-temporal representations in the deep feature space, such as OC-SVM and OC-NN (Ionescu et al., 2019; Sun et al., 2019). The drawback of such methods is that the trained models cannot be incrementally compatible with new data, leading to the need to retrain classifiers from scratch when new data is generated, such as in scenarios involving scene transitions. Another type of distance-based method is to use Gaussian mixture models (Wu et al., 2010) to model video feature normal vectors and measure deviations using Mahalanobis distance. In contrast, probability-based methods attempt to map the spatial-temporal representations of regular events into a probability framework and discriminate anomaly instances by measuring differences in probability distributions. Such methods tend to use traditional models such as Markov random fields (Benezeth et al., 2009) to build probability space. Deep learning-based attempts have encountered significant increases in computational costs and are noticeably slower in inference speeds.
@@
In fact, the most prevailing deep UVAD approach is reconstruction-based (Ramachandra et al., 2020). On the one hand, reconstruction-based methods (Liu et al., 2018b; Park et al., 2020; Liu et al., 2024b) aim to train models to represent the general spatial-temporal patterns of regular events through self-supervised proxy tasks, benefiting from advances in video self-supervised learning (Jing and Tian, 2020) and deep neural networks. On the other hand, such methods avoid complex mathematical computations, which are easy to implement and exhibit excellent performance, thus being widely praised by existing researchers. The premise of reconstruction-based methods is that generative models trained on massive normal samples can effectively infer the spatial-temporal patterns of regular events. For anomalous events, the performance of proxy tasks will significantly decrease, and the resulting error can be used as a quantitative basis for measuring deviations to calculate anomaly scores. Common generative models include deep autoencoders (Hasan et al., 2016; Chong and Tay, 2017; Liu et al., 2022a, c; Park et al., 2020), variational autoencoders (Fan et al., 2020; Yan et al., 2018), and generative adversarial networks (Chen et al., 2021; Huang et al., 2021; Nguyen et al., 2020; Li et al., 2022; Cai et al., 2021a; Zhang et al., 2022). Proxy tasks include reconstructing input sequences and predicting future frames, which belong to pixel-level image generation. However, most reconstruction-based methods only calculate frame-level errors as anomaly scores without performing spatial localization, as the spatial contribution of anomalies is typically not significant. Spatial localization can only serve as a quantitative visualization result.
@@ 3.1. Taxonomy and Advances
This article also systematically reviews deep learning-driven UVAD methods, providing a taxonomy that can aid in understanding the current state of research and inspire further exploration. However, due to space constraints and the emphasis of the tutorial paper on guiding beginners, we only present the underlying logic of the proposed taxonomy and summarize the research trends of the latest advances in the main text message. For a more detailed description of the existing UVAD methods, please refer to Section 1.1 of Supplement.
@@
The methods based on distance and probability described earlier focus more on handcrafted features (Cheng et al., 2015; Shi et al., 2021) and traditional classification (Vu et al., 2017) models, which have been primarily surpassed by distance-based methods in the era of deep learning. Therefore, existing classification systems appear outdated in delineating recent advancements and reflecting research trends, failing to highlight the latest challenges and directions. To address this, inspired by data preprocessing techniques and forms of deep learning modeling, we categorize UVAD into two main classes: Global Normality Learning (GNL) and Local Prototype Modeling (LPM).
@@
GNL utilizes the entire video sequence as input, often requiring no additional preprocessing such as spatial-temporal cube partitioning or foreground object extraction, and employs end-to-end deep neural networks to directly learn video ecologies (Huang et al., 2022; Gong et al., 2019; Park et al., 2020; Chang et al., 2020; Liu et al., 2024b, 2023e). Over the years, researchers have believed that videos possess two informational dimensions, namely spatial and temporal, corresponding to appearance and motion, requiring different approaches to capture their normality. Therefore, within our UVAD taxonomy, GNL is further subdivided into Single-Proxy Task (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020) and Multi-Proxy Task methods (Chang et al., 2020; Liu et al., 2022c, 2024b). In contrast, LPM methods argue that video data contain a plethora of redundant information tantamount to clues related to normality. Thus, they opt to use spatial-temporal cubes (Mahadevan et al., 2010; Roshtkhari and Levine, 2013; Sabokrou et al., 2017; Liu et al., 2021a) or foreground objects (Bao et al., 2022; Liu et al., 2023a) containing dense and effective information as network inputs instead of the entire video sequence, focusing on prototype feature learning of local patch. We classify such methods into spatial-temporal Patch-based and Object-driven methods.
UVAD methods can be grouped into single-task, multi-task, spatial-temporal patch-based, and foreground object-driven approaches. The first two categories, which input either full RGB frames or optical flow sequences, are considered global normality learning methods. In contrast, the latter two approaches focus on modeling local image patches or salient foreground objects, which are categorized under local prototype modeling. A taxonomy of these methods is presented in Table 2, showcasing their latest developments and interrelations. Specifically, single-task methods treat the spatial and temporal patterns in video as entangled, typically employing a single network structure to execute a unified task for learning spatiotemporal normality. These methods are easy to design and train but may perform suboptimally in handling diverse anomalies in complex scenes. In comparison, multi-task methods regard appearance and motion as distinct information dimensions, using multi-branch networks (e.g., parallel autoencoders or encoder-decoder architectures) to perform different tasks for learning spatial and motion normality separately. This approach effectively handles anomalies involving appearance, motion, or a combination of both and has demonstrated outstanding performance in industrial, traffic, and medical applications. LPM methods address the redundant information in raw image sequences, which increases data processing costs and introduces noise that can degrade model performance. These methods first identify relevant spatiotemporal regions through preprocessing techniques before modeling them. Specifically, spatial-temporal patch (STP) methods assume that anomalies occupy small spatiotemporal regions in the video, and thus, modeling local cubes of data allows for precise spatiotemporal anomaly localization. Foreground object-driven (FOD) methods focus on analyzing patterns of foreground objects, leveraging pre-trained object detection models to extract the regions of interest for subsequent modeling.
@@ 3.1.1. Global Normality Learning
Convolutional neural network-driven deep representation models can directly learn task-relevant spatial-temporal representations from raw video sequences and can adapt to different scale inputs and feature dimension requirements through simple structural adjustments. Global normality learning aims to learn video normality directly from complete RGB videos or optical flow sequences (Liu et al., 2023b). Compared to local prototype modeling, GNL requires no additional data preparation and is easy to optimize.
The earliest methods did not distinguish between spatial and temporal information, typically using only RGB videos as input and employing a single self-supervised proxy task (e.g., sequence reconstruction or future frame prediction (Zhao et al., 2024)) to broadly learn spatial-temporal normality. Generally, methods based on a single proxy task focus on designing more efficient single-stream end-to-end deep structures. Recent efforts include the introduction of 3D convolutional networks (Gong et al., 2019) and convolutional long short-term memory networks (Luo et al., 2017) to enhance the representation capability of spatial-temporal features. Subsequent researchers discovered that spatial and temporal information have different characteristics, with the former focusing on local pixel inference and the latter on modeling global dynamics. Moreover, the addition of proxy tasks such as reconstruction and prediction losses often brings additional performance gains without significantly increasing training costs. Thus, they proposed introducing additional proxy tasks within the GNL framework. In addition to the separation of spatial-temporal normality learning (Liu et al., 2022c; Chang et al., 2021) and simple proxy task stacking (Zhao et al., 2017; Chen et al., 2020), recent work has also explored novel tasks such as appearance-motion consistency (Cai et al., 2021b), spatial-temporal coherence (Cheng et al., 2023a; Ning et al., 2024), and correlation (Zhao et al., 2022).